OpenStudy (anonymous):

Find the slope between points (-3, 5) and (-3, 8). A -8/5 B 8/5 C 0 D undefined slope

OpenStudy (anonymous):

slope is (y2-y1)/(x2-x1) = (8-5)/(-3--3) = 3/0 This would be undefined because you cannot divide a number by zero and get an integer.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Anytime ;)

OpenStudy (anonymous):

more questions to come! look out! :)

OpenStudy (anonymous):

Haha shoot me a message and I'll do my best ;)